The supplement of an angle is equal to six times the complement of the angle. Find the size of the angle.

Answer by longjonsilver(2297)

let angle = x

supplement is 180-x
compliment is 90-x

180 - x = 6(90-x)
180 - x = 540 - 6x
180 + 5x = 540
5x = 360
x = 72

jon.